# Optimal Parallel Algorithms for Convex Hulls in 2D and 3D under Noisy Primitive Operations

### v1

@@ 1 Introduction
Work on searching and sorting 1-dimensional data with faulty comparisons originates with Rényi [29], who studied a binary search problem in 1961 where comparisons between two values return the wrong answer independently with probability $p<1/2$; see also, e.g., Pelc [25, 26]. In 1994, Feige, Raghavan, Peleg, and Upfal [13] showed how to sort $n$ items in $O(\log n)$ span with $O(n\log n)$ work in the CRCW PRAM model, with high probability,111 In this paper, we take “with high probability” (w.h.p.) to mean that the failure probability is at most $1/n^{c}$, for some constant $c\geq 1$. in the same noisy comparison model, and Leighton, Ma, and Plaxton [22] show how to achieve these bounds in the EREW PRAM model (we review the PRAM model of computation in Appendix A). Both algorithms make extensive use of the AKS sorting network [2, 3], however, which implies that their asymptotic performance bounds have very large constant factors. Recently, in a SPAA 2023 paper, Goodrich and Jacob [16] provide an efficient parallel sorting algorithm where comparisons are either persistently or independently faulty without using the AKS sorting network. Since computations cannot be fully correct with persistent errors, we focus in this paper on the design of efficient parallel computational geometry algorithms using Boolean geometric primitives, such as orientation queries or sidedness tests, that randomly return the wrong answer independently of previous queries with probability at most $p<1/2$ (where $p$ is known).
@@ 3 Failure Sweeping with Noisy Primitives
Failure sweeping, formalized by Ghouse and Goodrich [14], is a general technique used to improve confidence bounds on algorithms that either fail or exceed a desired span or work bound with some failure probability $p(n)$, given input size $n$. Say that our desired bounds for span and work done at a subproblem of size $n$ are $T(n)$ and $W(n)$. Our goal is for the entire algorithm to fail with probability at most $p(n)$, but when we recurse to sufficiently small subproblems of size $m$, e.g., $m=O(\log n)$, $p(m)$ may be exponentially larger than $p(n)$. If unchecked, errors in small subproblems propagate to the returned answer, meaning the algorithm can only guarantee success probability $1-p(m)$, rather than the intended $1-p(n)$. This is exactly the case in the noisy primitives setting. Algorithms resilient to noise run in span $T(m)$ if we wish for them to fail with probability at most $p(m)$. Attempting to have all subproblems fail with probability at most $p(n)$ would cause each subproblem to take span proportional to $T(n)$, leading to an inefficient algorithm. Failure sweeping gives us a way to improve this confidence bound from $1-p(m)$ to $1-p(n)$ without incurring this span blow-up.
@@
To complete the argument, one must show that not too many subproblems fail with probability $p(n)$. Specifically, if each subproblem of size $m$ fails with probability $p(m)$, we must show that no more than $q$ subproblems, determined in the analysis, fail with probability $1-p(n)$. Then, we apply our brute-force algorithm to compute the $q$ subproblems in span $T(n)$ and total work $\leq W(n)$. This way, all subproblems succeed with probability $1-p(n)$. Applied inductively, this shows that we can “upgrade” our failure probabilities at each level of recursion without asymptotically increasing span or work. This ultimately allows the algorithm to achieve confidence bound $1-p(n_{0})$, where $n_{0}$ is the initial problem size.
One challenge in designing failure sweep algorithms is in ensuring that $m$ is not too small and not too large compared to $n$. If $m$ is too small, then many subproblems are likely to fail, forcing us to recompute too many subproblems. If $m$ is too large, then the cost to recompute even a few subproblems may exceed $W(n)$. Indeed, this second issue arises in our 3D convex hull algorithm. We introduce a novel technique called size reduction, which is a preprocessing step that transforms large subproblems of size $m$ into $O(m)$ sufficiently small problems of total size $O(m)$ with probability $1-p(n)$. We are then able to show that not too many of these smaller subproblems fail, guaranteeing that our total brute-force work is no more than $W(n)$. How small these subproblems must be depends both on $W(n)$, the brute-force algorithm, and $p(m)$.
In the case where $p(n)$ is a high-probability bound in $n$, size reduction may be applied to any algorithm wherein this issue arises as long as one can efficiently transform a large subproblem into $O(m^{\alpha})$ sufficiently small subproblems of total size $O(m)$ with probability $1-p(n)$ for any fixed constant $\alpha$. Higher confidence guarantees may tolerate a larger fanout.
@@ 4 A Deterministic 2D Convex Hull Algorithm for CREW PRAM
We can then use $T_{j},R_{j},L_{j},q,p,$ and $r$ to determine whether $p$ is on the upper hull of $S_{j}$. With a constant number of orientation tests, we can determine whether either sequence of vertices $[L_{j},q,p,r,T_{j},R_{j}]$ or $[T_{j},q,p,r,R_{j},L_{j}]$ is a convex chain. Using the trivial repetition strategy, this takes $O(\log(m^{2}))$ time to guarantee with failure probability at most $m^{-2c}$. See Figure 2 for a visual of the following lemma.
@@
If either $[L_{j},q,p,r,T_{j},R_{j}]$ or $[T_{j},q,p,r,R_{j},L_{j}]$ is a convex chain, then $p$ is a member of the upper hull of $S_{j}$.
@@ 5 A Randomized 3D Convex Hull Algorithm for CREW PRAM
Attempting to apply the strategy that bounds the number of failed subproblems that we used for 2D convex hulls presents a problem. Recall from Reif and Sen’s original work [27] as well as Clarkson [8], if the problem at recursive level $l-1$ is size $M$, then its child problems at recursive level $l$ could be size at most $O(M^{7/8}\log M)$. Indeed, if we were to apply Lemma 5.1 to solve just a single large subproblem by brute-force, it would take $O(M^{7/2}\log M)$ work to solve it w.h.p. in $M$. In this section, we develop a novel technique we call size reduction to decompose these larger subproblems into problems that are small enough to recompute if needed but are large enough to have similar failure probabilities.
First, however, we observe that sufficiently small subproblems can be trivially recomputed via Lemma 5.1.
@@
We can recompute the values of all subproblems of size $m_{s}=O(M^{1/8})$ in $O(M\log M)$ work and $O(\log M)$ span w.h.p. in $M$.
Proof 5.4.
Recall that we choose a sample of size $|R|\leq M^{1/8}$ from which we constructed our cones. This implies that there are $O(M^{1/8})$ cones in our sampled hull, each of which induces one subproblem. The total size of all subproblems of size $O(M^{1/8})$ is at most $O(M^{1/8})\times O(M^{1/8})=O(M^{1/4})$. Applying the naive brute-force strategy of Lemma 5.1 gives us the required bound.
To illustrate our approach to size reduction, consider the following algorithm to compute a halfspace intersection of $m$ halfspaces w.h.p. in $M$. We can perform the original algorithm with the following two modifications. (1) Each recursive level must succeed with high probability in $M$, so no matter what the current problem size is we tune our noisy operations and sampling probabilities to fail with probability $\leq 1/M$. (2) Irrespective of the current problem size, always choose a sample of size $O(m^{1/8})$.
Each recursive level now takes span $O(\log M)$, total work $O(m\log M)$, and uses $O(m)$ noisy operations. However, because our samples are larger, we reduce the maximum subproblem size by roughly $O(m^{1/8})$ at each level, meaning that it takes at most eight levels of recursion such that the largest problem size is $O(m^{1/8})$. When we get to this level, we simply apply Lemma 5.1 to solve them w.h.p. in $M$ by brute-force. There are at most $O(m^{1/8})^{8}=O(m)$ such subproblems, so it takes $O(\log M)$ span and $O(m^{3/2}\log M)$ work to do this last step. Because all operations are done w.h.p. in $M$, there is no need to failure sweep these solutions. We simply recombine them as we return from the eight levels of recursion in an additional $O(\log M)$ span and $O(m\log M)$ work per level.
Lemma 5.5.
The intersection of a set of $m$ halfspaces can be computed in $O(m^{3/2}\log M)$ work and $O(\log M)$ span with high probability in $M$. This takes $O(m^{3/2})$ noisy operations.
We can modify this algorithm to instead operate as a preprocessing step prior to our initial recursion. For every subproblem of size $m=\omega(M^{1/8})$, perform the above steps for at most eight levels of recursion until we have decomposed $m$ into at most $O(m)$ subproblems of size $O(M^{1/8})$. Then we simply recurse on these induced subproblems, tuning parameters such that they each succeed high probability in $M^{1/8}$.
Recall that there are at most $O(M^{1/8})$ original subproblems of $M$. In total, we induce at most $O(M^{1/8}\times M)=O(M^{9/8})$ more subproblems.
Lemma 5.6.
Say that we have $O(M^{9/8})$ processes, each of which fail with high probability in $M^{1/8}$. Then at most a constant number of them fail with high probability in $M$.
Proof 5.7.
See Appendix E.3.
@@
Proof 5.9.
Follows from Theorem 5.2, Lemma 5.6, and an application of Lemma 5.5 to any failed subproblems. We note that the added subproblems induced by size reduction does no affect the result of Theorem 5.2, because total problem size over all subproblems is still $O(M)$ by pruning. The problems at the lowest level of this subproblem are of size at most $O(M^{1/8})$. Then, the cost to recompute a constant number of such subproblems with high-probability in $M$ is $O(\log M)$ span and $O((M^{1/8})^{3/2}\log M)\subset O(M\log M)$ work.
We observe that these are loose worst-case bounds on the number of problems and their size. In reality, pruning guarantees that each level of our recurrence, even the ones that are a part of size reduction, always sum up to $O(n)$. This implies that our induced subproblems all sum to $O(m)$ for each large subproblem of $M$. Therefore, even if we increase the number of subproblems that we recurse into, our total work does not increase because total work is proportional to problem size.
@@ Appendix C Review of Non-noisy Algorithms
We adapt our algorithm mostly from the optimal $O(\log n)$-span and $O(n\log n)$-work randomized 3D convex hull algorithm of Reif and Sen [27]. First, we note that Reif and Sen solve the dual problem of halfspace-intersection of $n$ halfspaces in 3D. This allows them to take advantage of an earlier result of Clarkson [8], who showed that a random sample of halfspaces $R$ can be used to divide the remaining $H\setminus R$ halfspaces into subproblems of roughly equal size, admitting a natural divide-and-conquer algorithm. Specifically, these good samples divide the remaining halfspaces into groups with the following two properties: (1) the largest group has size at most $O(|H|\log|R|/|R|)$ and (2) the sum over all groups is of size $O(|H|)$. However, Clarkson only showed that this is the case with constant probability $>1/2$.
@@ D.1 Omitted Proofs
Consider a subproblem $S^{\prime}$ of size $m^{2}$. To complete the proof, we will show that it can be solved in $O(\log(m^{2}))$ span and $O(m^{2}\log(m^{2}))$ work with failure probability no more than $m^{-2c^{\prime}}$. $S^{\prime}$ must combine $m$ subproblems of size $m$. We expect $m/m^{c^{\prime}}$ to fail by our inductive hypothesis, yet we require that no more than $m^{1/3}$ fail by Corollary 1. Let $X$ be a random variable representing the number of subproblems that fail. Markov’s inequality states that $\Pr[X\geq m^{1/3}]\leq(m/m^{c^{\prime}})\times(1/m^{1/3})=m^{2/3-c^{\prime}}$. To be a high probability bound in $m^{2}$, $22/3-c^{\prime}<-$ 2, so $c^{\prime}>8/3$.
@@ D.2 Additional Details
Figure 2: We depict an instance of the failure sweep brute-force algorithm determining that a point, $p$, is a part of the upper hull in $S_{j}$. In this case, $[L_{j},q,p,r,T_{j},R_{j}]$ is convex. We have depicted the lines $\overrightarrow{pr}$ and $\overrightarrow{pq}$ as well as the vertical line through $p$ that borders both regions defined by the lines. Notice, as the proof describes, no points in $S_{j}$ can exist in the region either to the left of $\overrightarrow{pr}$ and to the right of $\overrightarrow{pq}$, depicted with shading. It is clear that $p$ is an extreme point in $S_{j}$.
@@
If either $[L_{j},q,p,r,T_{j},R_{j}]$ or $[T_{j},q,p,r,R_{j},L_{j}]$ is a convex chain, then $p$ is a member of the upper hull of $S_{j}$.
@@
In either of the two cases, imagine we rotate the coordinate system such that $p$ has the highest $y$-coordinate of all $\{T_{j},R_{j},L_{j},p,q,r\}$. By the fact that no points exist to the left of $\overrightarrow{pr}$ and to the right of $\overrightarrow{pq}$, $p$ has the highest $y$-coordinate of all points in $S_{j}$ in this new coordinate system. Because the structure of a convex hull is invariant to rotation, $p$ must be on the convex hull of $S_{j}$. Returning to the original coordinate system, because either $[L_{j},p,T_{j}]$ or $[T_{j},p,R_{j}]$ is a right-hand turn and the fact that $L_{j}$, $T_{j}$, and $R_{j}$ are the leftmost, topmost, and rightmost points, $p$ must be on the upper hull of $S_{j}$, not its lower hull.
@@
Say that $m^{1/b}$ subproblems fail. To recompute them in parallel takes $O((m^{1/b}\times m)^{3/2}\times\log(m^{2}))$ $=O(m^{3/(2b)+3/2}\log(m^{2}))$ work. We can only afford $O(m^{2}\log(m^{2}))$ work at subproblem $S^{\prime}$, so $3/2b+3/2\leq 2$. We have that $b\geq 3$, so at most $O(m^{1/3})$ subproblems can fail such that $S^{\prime}$ can fix them in $O(m^{2}\log(m^{2}))$ work via the brute-force method described in the previous section.

### v2

@@ 1 Introduction
Work on searching and sorting 1-dimensional data with faulty comparisons originates with Rényi [29], who studied a binary search problem in 1961 where comparisons between two values return the wrong answer independently with probability $p<1/2$; see also, e.g., Pelc [25, 26]. In 1994, Feige, Raghavan, Peleg, and Upfal [13] showed how to sort $n$ items in $O(\log n)$ span with $O(n\log n)$ work in the CRCW PRAM model, with high probability, [^1] In this paper, we take “with high probability” (w.h.p.) to mean that the failure probability is at most $1/n^{c}$, for some constant $c\geq 1$. in the same noisy comparison model, and Leighton, Ma, and Plaxton [22] show how to achieve these bounds in the EREW PRAM model (we review the PRAM model of computation in Appendix A). Both algorithms make extensive use of the AKS sorting network [2, 3], however, which implies that their asymptotic performance bounds have very large constant factors. Recently, in a SPAA 2023 paper, Goodrich and Jacob [15] provide an efficient parallel sorting algorithm where comparisons are either persistently or independently faulty without using the AKS sorting network. Since computations cannot be fully correct with persistent errors, we focus in this paper on the design of efficient parallel computational geometry algorithms using Boolean geometric primitives, such as orientation queries or sidedness tests, that randomly return the wrong answer independently of previous queries with probability at most $p<1/2$ (where $p$ is known).
@@ 3 Failure Sweeping with Noisy Primitives
Failure sweeping, formalized by Ghouse and Goodrich [14], is a general technique used to improve confidence bounds on algorithms that either fail or exceed a desired span or work bound with some failure probability $p(n)$, given input size $n$. Say that our desired bounds for span and work done at a subproblem of size $n$ are $T(n)$ and $W(n)$. Our goal is for the entire algorithm to fail with probability at most $p(n)$, but when we recurse to sufficiently small subproblems of size $m$, e.g., $m=O(\log n)$, $p(m)$ may be exponentially larger than $p(n)$. If unchecked, errors in small subproblems propagate to the returned answer, meaning the algorithm can only guarantee success probability 1p(m)1-p(m), rather than the intended 1p(n)1-p(n). This is exactly the case in the noisy primitives setting. Algorithms resilient to noise run in span $T(m)$ if we wish for them to fail with probability at most $p(m)$. Attempting to have all subproblems fail with probability at most $p(n)$ would cause each subproblem to take span proportional to $T(n)$, leading to an inefficient algorithm. Failure sweeping gives us a way to improve this confidence bound from 1p(m)1-p(m) to 1p(n)1-p(n) without incurring this span blow-up.
@@
To complete the argument, one must show that not too many subproblems fail with probability $p(n)$. Specifically, if each subproblem of size $m$ fails with probability $p(m)$, we must show that no more than $q$ subproblems, determined in the analysis, fail with probability 1p(n)1-p(n). Then, we apply our brute-force algorithm to compute the $q$ subproblems in span $T(n)$ and total work $\leq W(n)$. This way, all subproblems succeed with probability 1p(n)1-p(n). Applied inductively, this shows that we can “upgrade” our failure probabilities at each level of recursion without asymptotically increasing span or work. This ultimately allows the algorithm to achieve confidence bound $1-p(n_{0})$, where $n_{0}$ is the initial problem size.
One challenge in designing failure sweep algorithms is in ensuring that $m$ is not too small and not too large compared to $n$. If $m$ is too small, then many subproblems are likely to fail, forcing us to recompute too many subproblems. If $m$ is too large, then the cost to recompute even a few subproblems may exceed $W(n)$. Indeed, this second issue arises in our 3D convex hull algorithm. We introduce a novel technique called size reduction, which is a preprocessing step that transforms large subproblems of size $m$ into $O(m)$ sufficiently small problems of total size $O(m)$ with probability 1p(n)1-p(n). We are then able to show that not too many of these smaller subproblems fail, guaranteeing that our total brute-force work is no more than $W(n)$. How small these subproblems must be depends both on $W(n)$, the brute-force algorithm, and $p(m)$.
In the case where $p(n)$ is a high-probability bound in $n$, size reduction may be applied to any algorithm wherein this issue arises as long as one can efficiently transform a large subproblem into $O(m^{\alpha})$ sufficiently small subproblems of total size $O(m)$ with probability 1p(n)1-p(n) for any fixed constant $\alpha$. Higher confidence guarantees may tolerate a larger fanout.
@@ 4 A Deterministic 2D Convex Hull Algorithm for CREW PRAM
We can then use $T_{j},R_{j},L_{j},q,p,$ and $r$ to determine whether $p$ is on the upper hull of $S_{j}$. With a constant number of orientation tests, we can determine whether either sequence of vertices $L_{j},q,p,r,T_{j},R_{j}]$ or $T_{j},q,p,r,R_{j},L_{j}]$ is a convex chain. Using the trivial repetition strategy, this takes $O(\log(m^{2}))$ time to guarantee with failure probability at most $m^{-2c}$. See Figure 2 for a visual of the following lemma.
@@
If either $L_{j},q,p,r,T_{j},R_{j}]$ or $T_{j},q,p,r,R_{j},L_{j}]$ is a convex chain, then $p$ is a member of the upper hull of $S_{j}$.
@@ 5 A Randomized 3D Convex Hull Algorithm for CREW PRAM
Attempting to apply the strategy that bounds the number of failed subproblems that we used for 2D convex hulls presents a problem. Recall from Reif and Sen’s original work [27] as well as Clarkson [8], if the problem at recursive level l1l-1 is size $M$, then its child problems at recursive level $l$ could be size at most $O(M^{7/8}\log M)$. Indeed, if we were to apply Lemma 5.1 to solve just a single large subproblem by brute-force, it would take $O(M^{7/2}\log M)$ work to solve it w.h.p. in $M$. In this section, we develop a novel technique we call size reduction to decompose these larger subproblems into problems that are small enough to recompute if needed but are large enough to have similar failure probabilities.
To illustrate our approach to size reduction, consider the following algorithm to compute a halfspace intersection of $m$ halfspaces w.h.p. in $M$. We can perform the original algorithm with the following two modifications. (1) Each recursive level must succeed with high probability in $M$, so no matter what the current problem size is we tune our noisy operations and sampling probabilities to fail with probability $\leq 1/M$. (2) Irrespective of the current problem size, always choose a sample of size $O(m^{1/16})$.
By Reif and Sen’s original work [27] as well as an analysis of Clarkson [8], each subproblem of size $c$ now takes $O(\log M)$ span and $O(\sqrt{m}\log M+c\log M)$ work to preprocess it into recursive subproblems. However, because our samples are larger, we reduce the maximum subproblem size by roughly $O(m^{1/16})$ at each level, meaning that it takes at most fifteen levels of recursion such that the largest problem size is $O(m^{1/8})$. When we get to this level, we simply apply Lemma 5.1 to solve them w.h.p. in $M$ by brute-force.
There can be at most $O(m)$ such subproblems so, in total, the cost of preprocessing and the cost of the final brute-force computation is at most $O(\log M)$ span and $O(m^{3/2}\log M)$ work. All operations are done w.h.p. in $M$, so failure sweeping is unnecessary. We simply recombine them as we return from the fifteen levels of recursion in an additional $O(\log M)$ span and $O(m\log M)$ work.
@@
The intersection of a set of $m$ halfspaces can be computed in $O(m^{3/2}\log M)$ work and $O(\log M)$ span with high probability in $M$.
We can modify this algorithm to instead operate as a preprocessing step prior to recursion. For every subproblem of size $m=\omega(\sqrt{M})$, perform the above steps for at most nine levels until we have decomposed $m$ into at most $O(M^{9/16})$ subproblems of size $O(\sqrt{M})$ (by pruning, the sum of subproblem sizes is no more than $O(M)$). This can be performed in $O(\log M)$ span and $O(M\log M)$ work. Then we recurse on all induced subproblems, tuning parameters such that each succeeds with high probability in $M^{1/2}$.
If each subproblem is of size $O(M^{1/2})$ and we can recompute a given subproblem of size $m$ by brute force in time $O(m^{3/2}\log M)$, then at most $O(M^{1/4})$ can fail such that we can recompute them all in $O(\log M)$ span and $O(M\log M)$ work. We note that, due to Theorem E.10, pruning, and the structure of the recursion, even if a subproblem is of size $o(M^{1/2})$, we can still tune parameters to evaluate it with high probability in $M^{1/2}$ without affecting total asymptotic complexity.
@@
Proof 5.5.
Follows from Theorem 5.2, our analysis above, and an application of Lemma 5.3 to any failed subproblems. We will now show that not too many fail w.h.p. in $M$. Let $X$ be a random variable representing the number of subproblems that fail and say that each subproblem fails with probability $M^{-c^{\prime}/2}$ for some constant $c^{\prime}$ that we will determine. Then $E[X]=O(M^{9/16})\times M^{-c^{\prime}/2}=O(M^{(9-8c^{\prime})/16})$. Recall that we require no more than $O(M^{1/4})$ subproblems to fail. Applying Markov’s inequality, we have that
$\Pr[X\geq M^{1/4}]=O(M^{(9-8c^{\prime})/16})\times M^{-1/4}$
$O(M^{(5-8c^{\prime})/16}).$
To be a high-probability bound, this failure probability must be less than $M^{-1}$. Then we can recompute all failed subproblems w.h.p. in $M$ so long as $(5-8c^{\prime})/16<-1$, and so $c^{\prime}>21/8$.
@@ Appendix C Review of Non-noisy Algorithms
We adapt our algorithm mostly from the optimal $O(\log n)$-span and $O(n\log n)$-work randomized 3D convex hull algorithm of Reif and Sen [27]. First, we note that Reif and Sen solve the dual problem of halfspace-intersection of $n$ halfspaces in 3D. This allows them to take advantage of an earlier result of Clarkson [8], who showed that a random sample of halfspaces $R$ can be used to divide the remaining $H\setminus R$ halfspaces into subproblems of roughly equal size, admitting a natural divide-and-conquer algorithm. Specifically, these good samples divide the remaining halfspaces into groups with the following two properties: (1) the largest group has size at most $O(|H|\log|R|/|R|)$ and (2) the sum over all groups is of size $O(|H|)$. However, Clarkson only showed that this is the case with constant probability >1/2>1/2.
@@ D.1 Omitted Proofs
Consider a subproblem $S^{\prime}$ of size $m^{2}$. To complete the proof, we will show that it can be solved in $O(\log(m^{2}))$ span and $O(m^{2}\log(m^{2}))$ work with failure probability no more than $m^{-2c^{\prime}}$. $S^{\prime}$ must combine $m$ subproblems of size $m$. We expect $m/m^{c^{\prime}}$ to fail by our inductive hypothesis, yet we require that no more than $m^{1/3}$ fail by Corollary 1. Let $X$ be a random variable representing the number of subproblems that fail. Markov’s inequality states that $\Pr[X\geq m^{1/3}]\leq(m/m^{c^{\prime}})\times(1/m^{1/3})=m^{2/3-c^{\prime}}$. To be a high probability bound in $m^{2}$, $2/3-c^{\prime}<-2$, so $c^{\prime}>8/3$.
@@ D.2 Additional Details
Figure 2: We depict an instance of the failure sweep brute-force algorithm determining that a point, $p$, is a part of the upper hull in $S_{j}$. In this case, $L_{j},q,p,r,T_{j},R_{j}]$ is convex. We have depicted the lines $\overrightarrow{pr}$ and $\overrightarrow{pq}$ as well as the vertical line through $p$ that borders both regions defined by the lines. Notice, as the proof describes, no points in $S_{j}$ can exist in the region either to the left of $\overrightarrow{pr}$ and to the right of $\overrightarrow{pq}$, depicted with shading. It is clear that $p$ is an extreme point in $S_{j}$.
@@
If either $L_{j},q,p,r,T_{j},R_{j}]$ or $T_{j},q,p,r,R_{j},L_{j}]$ is a convex chain, then $p$ is a member of the upper hull of $S_{j}$.
@@
In either of the two cases, imagine we rotate the coordinate system such that $p$ has the highest $y$-coordinate of all $\{T_{j},R_{j},L_{j},p,q,r\}$. By the fact that no points exist to the left of $\overrightarrow{pr}$ and to the right of $\overrightarrow{pq}$, $p$ has the highest $y$-coordinate of all points in $S_{j}$ in this new coordinate system. Because the structure of a convex hull is invariant to rotation, $p$ must be on the convex hull of $S_{j}$. Returning to the original coordinate system, because either $L_{j},p,T_{j}]$ or $T_{j},p,R_{j}]$ is a right-hand turn and the fact that $L_{j}$, $T_{j}$, and $R_{j}$ are the leftmost, topmost, and rightmost points, $p$ must be on the upper hull of $S_{j}$, not its lower hull.
@@
Say that $m^{1/b}$ subproblems fail. To recompute them in parallel takes $O((m^{1/b}\times m)^{3/2}\times\log(m^{2}))$ $O(m^{3/(2b)+3/2}\log(m^{2}))$ work. We can only afford $O(m^{2}\log(m^{2}))$ work at subproblem $S^{\prime}$, so $3/2b+3/2\leq 2$. We have that $b\geq 3$, so at most $O(m^{1/3})$ subproblems can fail such that $S^{\prime}$ can fix them in $O(m^{2}\log(m^{2}))$ work via the brute-force method described in the previous section.